Let R be a binary relation on the set of all positive integers such that R = { (a,b) | a- b is an odd positive integer } Is R reflexive, symmetric, antisymmetric, transitive? Is R an equivalence relation? A partial ordering relation
1
Expert's answer
2020-11-05T14:19:58-0500

(b) "R = \{ (a,b) | a-b \ \text{is an odd positive integer }\}"

  • Reflexivity. For all "x" we have "(x, x) \notin R" because "x-x=0" is not a positive integer. The realtion is NOT reflexive.

Symmetry. Let "(x, y) \in R". Then "x-y" is a positive integer and "y-x = -(x-y)" is negative. Therefore, "(y, x) \notin R" and the relation is NOT symmetric because the following is always true:

"(x, y) \in R \wedge (y, x) \in R \implies x=y"

Antisymmetry. Let "(x, y) \in R" and "(y, x) \in R". Then "x-y" is a positive integer and "y-x" is a positive integer from the definition. Therefore "x-y>0" and "y-x>0". There are no such numbers "x, y" that satisfy "x>y" and "y>x" . Therefore, "(x, y) \in R \wedge (y, x) \in R" is a contradiction and the relation is antisymmetric because the following is always true:

"(x, y) \in R \wedge (y, x) \in R \implies x=y"
  • Transitivity. Let "(x, y) \in R \wedge (y,z) \in R." and "x \subseteq y \wedge y \subseteq x". Then for all"a \in x" we have "a \in y", and therefore "a \in z". Thus "x \subseteq z"Then "x-y" is an odd positive integer and "y-z" is a odd positive integer from the definition. "x-z = (x-y)+(y-z) - 2y". The difference of two even numbers is even, therefore "(x,z) \notin R". The relation is NOT transitive.

Answer. R can be neither an equivalence relation nor a partial ordered set because it is at least not reflexive.